OpenStudy (anonymous):

turn x^2-12x+36 into this format (x +- x) (x +- x)

OpenStudy (anonymous):

pls help

OpenStudy (tkhunny):

First, that is horrible notation. The symbol 'x' is used in at least two different ways. Very confusing. This might be better: (x +/- a)(x +/- b) where a*b = 36

OpenStudy (tkhunny):

Also, a + b = -12 This gives us a clue. We can no refine the form: (x - a)(x - b) where a*b = 36 and a + b = -12 You must see why this is so before proceeding.

OpenStudy (anonymous):

ok i see

OpenStudy (anonymous):

i see where did you go

OpenStudy (anonymous):

explain better if you understand.

OpenStudy (anonymous):

it would be (x +- (a) multiple of 36) (x +- (b)multiple of 36)

OpenStudy (tkhunny):

It's fine the way it was. (x - a)(x - b) where a*b = 36 and a + b = -12 For a*b = 36, we have these possiblities: -36 * -1 -18 * -2 -12 * -3 -9 * -4 -6 * -6 36 * 1 18 * 2 12 * 3 9 * 4 6 * 6 For a + b = -12?? Can we find a pair that does that?